OpenStudy (anonymous):

if the measure of arc hi is 80 and the measure of arc ji is 75 calculate m hki answers- a 105 b 52.5 c 210 d 130

OpenStudy (ashleyisakitty):

= 90d-m<IHJ = 90d-(m<arc JI)/2 = 90d-75d/2 = 90d-37.5d=52.5d m<HKI=2m<HKM=2*52.5d =105d